OpenStudy (anonymous):

x + 5y = -45 -2x - 15y = 140 elimination method

OpenStudy (anonymous):

you can either eliminate x or y in this simultaneous equation. Let's try eliminating x. x+5y =-45 Multiply both sides by 2 2x+10y = -90 add this equation to -2x -15y = 140 we get -5y =50 simplify 5y = -50 y = -10 now substitute y into any equation to find x

OpenStudy (anonymous):